OpenStudy (tayyllllorr):

will award a medal At an amusement park, the probability that a child eats popcorn and cotton candy is 0.58. The probability that a child eats popcorn is 0.69, and the probability that a child eats cotton candy is 0.87. What is the probability (rounded to the nearest hundredth) that a child eats popcorn given that the child has already eaten cotton candy? 0.84 0.67 0.40 0.79

OpenStudy (sshayer):

P(A/B)\[P \left( A/B \right)=\frac{ P \left( A~ and ~ B \right) }{ P \left( B \right) }\]